OpenStudy (anonymous):

PLEASE HELP!!! What is the 4th term of (a-√2)^8 What is the 6th term of (2x-3y)^11 What is the 2nd term of (y-x)^4 PLEASE HELP!!!

OpenStudy (anonymous):

Are you familiar with the Binomial Theorem, or at least Pascal's Triangle?

OpenStudy (anonymous):

Yes to the binomial theorem but I have difficulty using it

OpenStudy (anonymous):

Expand them

OpenStudy (anonymous):

How though?

OpenStudy (anonymous):

\[(a+b)^n=\sum_{k=0}^n\begin{pmatrix}n\\k\end{pmatrix}a^{n-k}b^k\] The fourth term of \(\left(a-\sqrt2\right)^8\) is then \[\begin{pmatrix}8\\3\end{pmatrix}a^{8-3}\left(\sqrt2\right)^3=\frac{8!}{3!5!}\times 2^{3/2}a^5=\cdots\]

OpenStudy (anonymous):

Do the same for the rest. Keep track of the index: keep in mind that \(k=0\) gives you the first term, so \(k=1\) gives the second, and so on.

OpenStudy (anonymous):

Ok so would the final answer be -112√(2) a^5 ... (I have multiple choice)